OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2.

OpenStudy (anonymous):

Okay, the function I picked from number 2 is f(x) = 10x + 6 - 3. I do not understand what the heck I'm doing lol. The h(x) I decided to go with is h(x) = 2x + 6 and the x value i chose is x = 5. I'm not sure what to do next or how to set it up. So far, I tried 10(2(5)+6)+6-3 which feels completely wrong. lol, please help!

OpenStudy (perl):

you can pick any old function h(x) = x + 2 f(x) = x - 8

OpenStudy (anonymous):

I picked h(x) = 2x +6 The issue is not knowing what to do with it.